OpenStudy (anonymous):

d=rt+5 solve for r **show work**

jimthompson5910 (jim_thompson5910):

\[\Large d=rt+5\] \[\Large d-5=rt\] \[\Large \frac{d-5}{t}=r\] \[\Large r=\frac{d-5}{t}\]